# Large-Scale Training of Graph Neural Networks for Optimal Markov-Chain Partitioning Using the Kemeny Constant

### v2

@@ Abstract
The recent trend in using network and graph structures to represent a variety of different data types has renewed interest in the graph partitioning (GP) problem. This interest stems from the need for general methods which can both efficiently identify network communities and reduce the dimensionality of large graphs while satisfying various application-specific criteria. Traditional clustering algorithms often struggle to capture the complex relationships within graphs and generalise to arbitrary clustering criteria. The emergence of graph neural networks (GNNs) as a powerful framework for learning representations of graph data provides new approaches to solving the problem. Previous work has shown GNNs to be capable of proposing partitionings using a variety of criteria, however, these approaches have not yet been extended to work on Markov chains or kinetic networks. These arise frequently in the study of molecular systems and are of particular interest to the biochemical modelling community. In this work, we propose several GNN-based architectures to tackle the graph partitioning problem for Markov Chains described as kinetic networks. This approach aims to minimize how much a proposed partitioning changes the Kemeny constant. We propose using an encoder-decoder architecture and show how simple GraphSAGE-based GNNs with linear layers can outperform much larger and more expressive attention-based models in this context. As a proof of concept, we first demonstrate the method’s ability to cluster randomly connected graphs. We also use a linear chain architecture corresponding to a 1D free energy profile as our kinetic network. Subsequently, we demonstrate the effectiveness of our method through experiments on a data set derived from molecular dynamics. We compare the performance of our method to other partitioning techniques such as PCCA+. We explore the importance of feature and hyperparameter selection and propose a general strategy for large-scale parallel training of GNNs for discovering optimal graph partitionings.
@@ 1 Introduction
Markov chains (MCs), and their extensions, have become a key tool across multiple domains to model dynamical systems[1, 2]. In their simplest form, MCs incorporate sequential data as a Markovian process into a stochastic model, where the probability of observing the next event depends solely on the preceding event. By eliminating the long timescale dependencies that often occur in complex systems, MCs offer a robust and formal method to interpret high-dimensional datasets in the emergent era of big data. The simplicity, transferability, and capability of MCs reduce the difficulties associated with modelling and interpreting ordered dynamics, and have led to a wide range of use cases. These applications encompass modelling character sequences in computer science and genomics[3, 4], predicting web-user activity[5], and modelling time-dependent processes in physics and chemistry[6, 7], to name just a few.
@@
A field that has widely adopted these techniques is computational biochemistry, where the Markov State Model (MSM) formalism has emerged as a strategy for interpreting the results of molecular dynamics (MD) simulations as transitions between aggregated conformational states[8, 9, 10]. The surge in popularity of MSMs mirrors researchers’ increased ability to quickly generate large quantities of MD data, providing a useful tool to condense massive data sets into a small and intuitive probabilistic format. MSMs offer intuitive descriptions of complex dynamic systems as transitions between stable minima and unstable transition states[11], helping to elucidate the underlying functional kinetics of interest[12, 13]. Over the past few decades, research in this community has led to several domain-specific advances in developing techniques to automatically generate MSM representations of molecular systems. Simultaneously, these techniques have been employed to further sample the multitude of different configurations of biomolecular systems[14, 15].
@@
Despite being smaller than the huge data sets they originate from, many MSMs remain high-dimensional, consisting of potentially hundreds of states. This large number of states is often unmanageable for downstream tasks such as drug design and biomolecular engineering[16]. As size often becomes the limiting factor to establishing an intuitive understanding of systems, many different approaches have been proposed to reduce the size of MCs and generate optimal ones [17, 18, 19, 20]. However, our ability to generate large data sets is only increasing, as modern coarse-graining protocols[21] and new developments in machine learning[22] allow for larger time scales to be probed. Traditional graph partitioning algorithms, such as spectral clustering, k-means, or multi-level methods, and their derivatives, have been widely employed to address this problem [23, 24, 25]. However, these algorithms often face limitations in capturing the complex and intricate dependencies on global structure in MCs. Additionally, the generality of these methods means they often disregard the critical dynamic information held within MCs. Alternative Markov-Chain Monte Carlo (MCMC) approaches have been applied for dimensionality reduction on MCs, which can be used with most clustering criteria [26, 27, 28]. Nevertheless, these approaches encounter difficulties in efficiently sampling the most relevant regions of partitioning space, resulting in impractical computation times and frequently overlooking the most optimal solutions.
@@
In this paper, we build upon previous work [29] to create a general methodology for reducing the dimensionality of MCs, leveraging recent developments in Machine Learning (ML), where Graph Neural Networks (GNNs) are applied to solve the more general graph partitioning (GP) problem[30, 31, 32]. Although several existing approaches use GNNs to minimise partitioning cuts[33] or modularity[34], these applications do not extend to kinetic graphs or dynamical criteria. Additionally, these approaches mostly focus on providing approximate partitionings for large graphs rather than exploring their potential for identifying optimal partitionings in smaller networks. We extend these methods by representing MCs as kinetic networks and show how performing gradient descent on a large number of randomly initialised small GNNs can yield optimal network partitionings using elaborate optimization criteria. As a metric to quantitatively assess the degree to which a clustering preserves the underlying dynamics of the MSM, we use the change of Kemeny constant (KC). KC is an inherently dynamical quantity derived from the Mean First Passage Times (MFPTs) associated with any kinetic network, including those used within computational biochemistry[35]. The proposed MC clustering method can produce a faithful low-dimensional replication of a system’s underlying dynamics, while also providing a quantifiable measurement related to the amount of kinetic information lost. Along with outlining a general strategy for the large-scale training of GNNs for graph partitioning, we show our method’s performance when minimising the change in KC on several different kinds of synthetic graphs and an example derived from MD simulations. We also test more complex GNN architectures, showing how large-scale training of simple GNNs is the most effective for the task and hypothesizing why this is the case.
@@ 2.1 Graphs
A graph ${G}$ with $N$ nodes can be represented as a set of vertices ${V}=\{v_{1},...,v_{N}\}$ and edges $e_{ij}\in{E}$, which are often represented as a binary adjacency matrix, $\mathbf{A}$, where:
$A_{ij}=\begin{cases}1,&\text{if}\ e_{ij}\in E\\ 0,&otherwise\end{cases}$ (1)
Each node, $v_{i}$, can be associated with a $d_{\text{feat}}$ dimensional feature vector, $f_{i}$, similarly admitting a matrix representation as a $(N\times d)$ matrix, $\mathbf{F}$. These features can be derived from the underlying data or generated at run time in the case of a featureless graph. This makes it possible to define a general graph structure as:
@@
In what follows, the terms clustering, partitioning, and lumping are used interchangeably to refer to the GP problem. The general partitioning problem consists of dividing the nodes of $G$ into $M$ discrete non-overlapping subsets, $s_{m}\in S$, such that specific criteria are met. These objectives can be general and applicable to nearly all graphs, and include, e.g., optimising connectivity, minimising inter-cluster edges, or maximising intra-cluster similarities. Furthermore, these objectives can also be domain-specific, based on the underlying model the network represents, such as how the KC represents the dynamical information of the system. The $(N\times M)$ partitioning assignment matrix, $\mathbf{S}$, assigns the nodes of $G$ into $M$ distinct states, $\{s_{1},\ldots,s_{m}\}$, where:
$S_{ik}=\begin{cases}1,&\text{if}\ {v}_{i}\in s_{k}\\ 0,&otherwise\end{cases}$ (3)
Since the discrete nature of the assignment matrix poses difficulties for gradient-based methods, it is common to relax this constraint and work with continuous probabilities instead. In practice, this means introducing a soft assignment matrix instead, where $S_{ik}$ now denotes the probability of node $v_{i}$ belonging to state $s_{k}$, subject to the conservation of probability:
$\sum^{M}_{k=1}S_{ik}=1\ \forall\ v_{i}\in V$
where $k$ runs over all $M$ possible states.
Two of the most commonly used clustering criteria used for the GP problem are the modularity and the minimum cut, which are defined as follows:
Modularity: The modularity, $Q$, measures how the connectivity of a graph partitioning differs from what would be expected in a purely random graph with the same node degree distribution. Consequently, this metric provides a quantitative assessment of how statistically surprising a proposed partitioning is, with higher values implying the presence of underlying communities within the data [36]. The modularity is calculated using the equation:
$Q=\frac{1}{2m}\sum_{ij}^{N}\sum_{k}^{M}\bigg{[}A_{ij}-\frac{d_{i}d_{j}}{2m}\bigg{]}\ {S}_{ik}{S}_{jk}$ (4)
where $m$ is the total number of edges in the graph, $d_{i}$ and $d_{j}$ are the node degrees of nodes $v_{i}$ and $v_{j}$, respectively, $A_{ij}$ is given by Eq. (1), and $S_{ik}$ and $S_{jk}$ are given by Eq. (3). While there have been criticisms of the underlying assumptions for modularity, it is a widely used and useful partitioning criterion in many areas[37, 38, 39].
@@
$C=\sum^{N}_{i=1}\sum^{N}_{j\neq i}\sum^{M}_{k=1}\sum^{M}_{l\neq k}S_{ik}S_{jl}A_{ij}$ (5)
Where all indices and matrix values are given as previously defined. The simplicity of this criterion has led to its popularity across multiple domains, spurred on by its proved equivalence to the max-flow criterion [40]. It stands as one of the most studied partitioning criteria [41, 42, 43], alongside its further generalisations commonly used to obtain more balanced clusters with minimal cuts [44, 45].
@@ 2.2 Markov Chains
The dynamical clustering criterion based on KC we used in this paper is derived from the kinetic properties of an MC. In previous work, we explored the relationship between the MFPTs and KC, providing thorough derivations of all the quantities presented below. Here we briefly summarise the key results required and direct the reader to our previous paper for a detailed reference[46].
In a discrete-time, homogeneous MC comprising of $n$-states, the probability vector of a random walk occupying any of the $n$-states at time $t$, $\mathbf{p}(t)$, can be expressed using a stochastic rate matrix, $\mathbf{K}$, and a set of initial occupation probabilities, $\mathbf{p}(0)$, such that:
@@
Where $t=l\tau$, since in the discrete case moves between states happen at multiples $l=1,2,\ldots$ of a given time interval $\tau$. The quantity $\mathbf{Q}(t)=e^{\mathbf{K}\tau}$ is known as the Markov matrix, and to ensure the conservation of probability, we have:
@@
Where the index $l$ orders the eigenvalues in descending order.
This yields real eigenvalues that are upper-bounded by 0. The eigenvector corresponding to the largest right eigenvector, $\psi^{(1)}$, is known as the equilibrium probability and is commonly denoted by $\mathbf{p}^{\text{eq}}$, corresponding to the distribution the system converges to as $t\to\infty$. The other eigenvectors contain useful kinetic information corresponding to the relaxation and mixing times of the MC. These eigenvectors can be used to derive an expression for the average time it takes for a random walk starting from state $i$ to first arrive at state $j$, i.e., the MFPTs of an MC, denoted $t_{ji}$:
@@
This constant provides an inherently dynamical scalar quantity, ideal for gradient-based optimisation methods. By reintroducing the assignment matrix, $\mathbf{S}$, from Eq. (3), and defining a partitionings lumped equilibrium probability, $P^{\text{eq}}_{M}$, as simply:
$P^{\text{eq}}_{M}=\sum_{i\in M}p_{i}S_{iM}$ (12)
KC can be redefined in terms of a partitioned MC’s KC, $\zeta_{\textbf{S}}$, and a variational term, such that:
$\zeta=\zeta_{\textbf{S}}+\sum_{M}\frac{1}{P^{\text{eq}}_{M}}\sum_{i,j\in M}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}$ (13)
Rearranging this equation yields provides our partitioning metric $\Delta K$:
$\zeta-\zeta_{\textbf{S}}=\Delta K=\sum_{M}\frac{1}{P^{\text{eq}}_{M}}\sum_{i,j\in M}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}$ (14)
From this equation, it is clear that for the trivial single-state partitioning $M=1$, $\Delta K=\zeta$, and when $M=N$, the value of $\Delta K$ is 0.

### v3

# Kemeny Constant-Based Optimization of Network Clustering Using Graph Neural Networks

@@ Abstract
The recent trend in using network and graph structures to represent a variety of different data types has renewed interest in the graph partitioning (GP) problem. This interest stems from the need for general methods that can both efficiently identify network communities and reduce the dimensionality of large graphs while satisfying various application-specific criteria. Traditional clustering algorithms often struggle to capture the complex relationships within graphs and generalize to arbitrary clustering criteria. The emergence of graph neural networks (GNNs) as a powerful framework for learning representations of graph data provides new approaches to solving the problem. Previous work has shown GNNs to be capable of proposing partitionings using a variety of criteria. However, these approaches have not yet been extended to Markov chains or kinetic networks. These arise frequently in the study of molecular systems and are of particular interest to the biomolecular modeling community. In this work, we propose several GNN-based architectures to tackle the GP problem for Markov Chains described as kinetic networks. This approach aims to maximize the Kemeny constant, which is a variational quantity and it represents the sum of time scales of the system. We propose using an encoder-decoder architecture and show how simple GraphSAGE-based GNNs with linear layers can outperform much larger and more expressive attention-based models in this context. As a proof of concept, we first demonstrate the method’s ability to cluster randomly-connected graphs. We also use a linear chain architecture corresponding to a 1D free energy profile as our kinetic network. Subsequently, we demonstrate the effectiveness of our method through experiments on a data set derived from molecular dynamics. We compare the performance of our method to other partitioning techniques such as PCCA+. We explore the importance of feature and hyperparameter selection and propose a general strategy for large-scale parallel training of GNNs for discovering optimal graph partitionings.
@@ 1 Introduction
Markov chains (MCs), and their extensions, have become a key tool across multiple domains to model dynamical systems1, 2. In their simplest form, MCs incorporate sequential data as a Markovian process into a stochastic model, where the probability of observing the next event depends solely on the preceding event. By eliminating the long timescale dependencies that often occur in complex systems, MCs offer a robust and formal method to interpret high-dimensional data sets in the emergent era of big data. The simplicity, transferability, and capability of MCs reduce the difficulties associated with modelling and interpreting ordered dynamics, and have led to a wide range of use cases. These applications encompass modeling character sequences in computer science and genomics3, 4, predicting web-user activity5, and modelling time-dependent processes in physics and chemistry6, 7, to name just a few.
@@
An application that has widely adopted these techniques is biomolecular simulation, where the Markov state model (MSM) formalism has emerged as a strategy for interpreting the results of molecular dynamics (MD) simulations as transitions between aggregated conformational states8, 9, 10. The surge in popularity of MSMs mirrors the increased ability to quickly generate large quantities of MD data, providing a useful tool to condense massive data sets into a small and intuitive probabilistic format. MSMs offer intuitive descriptions of complex dynamic systems as transitions between metastable minima and saddle points serving as transition states11, helping to elucidate the underlying functional kinetics of interest12, 13. Over the past few decades, several domain-specific advances led to the development of techniques to automatically generate MSM representations of molecular systems. Simultaneously, these techniques have been employed to further sample the multitude of different configurations of biomolecular systems14, 15.
@@
Despite being smaller than the huge data sets from which they originate, many MSMs remain high-dimensional, potentially consisting of hundreds of states. This large number of states is often unmanageable for downstream tasks such as drug design and biomolecular engineering16. As size often becomes the limiting factor to establishing an intuitive understanding of systems, many different approaches have been proposed to reduce the size of MCs and generate optimal ones 17, 18, 19, 20. However, the ability to generate large data sets is only increasing, as modern coarse-graining protocols21 and new developments in machine learning22 allow for larger time scales to be probed. Traditional graph partitioning (GP) algorithms, such as spectral clustering, k-means, or multilevel methods, and their derivatives, have been widely employed to address this problem 23, 24, 25. However, these algorithms often face limitations in capturing the complex and intricate dependencies on global structure in MCs. Additionally, the generality of these methods means that they often disregard the critical dynamic information held within MCs. Alternative Markov-Chain Monte Carlo (MCMC) approaches have been applied for dimensionality reduction on MCs, which can be used with most clustering criteria 26, 27, 28. Nevertheless, these approaches encounter difficulties in efficiently sampling the most relevant regions of partitioning space, resulting in impractical computation times and frequent overlooking of the most optimal solutions.
@@
In this paper, we introduce machine learning (ML) models based on graph neural networks (GNN) to optimize dynamical network clusterings. We build upon previous work 29 for reducing the dimensionality of MCs, and propose ML architectures leveraging recent developments in using GNNs to solve the more general GP problem30, 31, 32. Although several existing approaches use GNNs to minimize partitioning cuts33 or modularity34, these applications do not extend to kinetic graphs or optimization based on dynamical criteria. Additionally, these approaches mostly focus on providing approximate partitionings for large graphs rather than exploring their potential for identifying optimal partitionings in smaller networks. We extend these methods by representing MCs as kinetic networks and show how performing gradient descent on a large number of randomly initialized small GNNs can yield optimal network partitionings using elaborate optimization criteria. As a metric to quantitatively assess the degree to which a clustering preserves the underlying dynamics of the MSM, we use the change of Kemeny constant (KC). KC is an inherently dynamical quantity derived from mean first passage times (MFPTs) associated with any kinetic network, including those used within biomolecular simulations35. We also compare our clustering results with a commonly used kinetic network clustering method for MSMs, the robust Perron-cluster analysis (PCCA+), which does not use the KC as the optimization criterion36.
The proposed MC clustering method can produce a faithful low-dimensional replication of a system’s underlying dynamics, while also providing a quantifiable measurement related to the amount of kinetic information lost. Along with outlining a general strategy for the large-scale training of GNNs for graph partitioning, we show our method’s performance when minimizing the change in KC on several different kinds of synthetic graphs and an example derived from MD simulations. We also test more complex GNN architectures, showing how large-scale training of simple GNNs is the most effective for the task and hypothesizing why this is the case.
@@ 2.1 Graphs
A graph ${G}$ with $N$ nodes can be described as a set of vertices ${V}=\{v_{1},...,v_{N}\}$ and edges $e_{ij}\in{E}$, which are often represented as a binary adjacency matrix, $\mathbf{A}$, where:
$A_{ij}=\begin{cases}1,&\text{if}\ e_{ij}\in E\\ 0,&\text{otherwise}\end{cases}$ (1)
Each node, $v_{i}$, can be associated with a $d_{\text{feat}}$ dimensional feature vector, $F_{i}$, similarly admitting a matrix representation as a $(N\times d_{\text{feat}})$ matrix, $\mathbf{F}$. These features can be derived from the underlying data or generated at run time in the case of a featureless graph. This makes it possible to define a general graph structure as:
@@
In what follows, the terms clustering and partitioning are used interchangeably to refer to the GP problem. The general partitioning problem consists of dividing the nodes of $G$ into $M$ discrete non-overlapping subsets, $L\in\{1,\ldots,M\}$, such that specific criteria are met. These objectives can be general and applicable to nearly all graphs, and include, e.g., optimizing connectivity, minimizing inter-cluster edges, or maximizing intra-cluster similarities. Furthermore, these objectives can also be domain-specific, based on the underlying model the network represents, such as how the KC represents the dynamical information of the system. The $(N\times M)$ partitioning assignment matrix, $\mathbf{S}$, assigns the nodes of $G$ into $M$ distinct states, such that:
$S_{iL}=\begin{cases}1,&\text{if}\ {v}_{i}\in L\\ 0,&\text{if}\ {v}_{i}\notin L\end{cases}$ (3)
Since the discrete nature of the assignment matrix poses difficulties for gradient-based methods, it is common to relax this constraint and work with continuous probabilities instead. In practice, this means introducing a soft assignment matrix instead, where $S_{iL}$ now denotes the probability of node $v_{i}$ belonging to cluster $L$, subject to the conservation of probability:
$\sum^{M}_{L}S_{iL}=1\ \forall\ v_{i}\in V$
where $L$ runs over all $M$ possible clusters.
Some of the most commonly used clustering criteria used for the GP problem are the modularity, minimum cut, and Davies-Bouldin index (DBI), which are defined as follows:
Modularity: The modularity, $Q$, measures how the connectivity of a graph partitioning differs from what would be expected in a purely random graph with the same node degree distribution. Consequently, this metric provides a quantitative assessment of how statistically surprising a proposed partitioning is, with higher values implying the presence of underlying communities within the data 37. The modularity is calculated using the equation:
$Q=\frac{1}{2m}\sum_{ij}^{N}\sum_{L}^{M}\bigg{[}A_{ij}-\frac{d_{i}d_{j}}{2m}\bigg{]}\ {S}_{iL}{S}_{jL}$ (4)
where $m$ is the total number of edges in the graph, $d_{i}$ and $d_{j}$ are the degrees of nodes $v_{i}$ and $v_{j}$, respectively, $A_{ij}$ is given by Eq. (1), and $S_{iL}$ and $S_{jL}$ are given by Eq. (3). Networks characterized by high modularity exhibit dense intra-cluster node connections, but sparse inter-cluster node connections. While there have been criticisms of the underlying assumptions for modularity, it is a widely used and useful partitioning criterion in many areas38, 39, 40.
@@
$C=\sum^{N}_{i}\sum^{N}_{j\neq i}\sum^{M}_{K}\sum^{M}_{L\neq K}S_{iK}S_{jL}A_{ij}$ (5)
where all indices and matrix values are given as previously defined. The simplicity of this criterion has led to its popularity across multiple domains, spurred on by its proved equivalence to the max-flow criterion 41. It stands as one of the most studied partitioning criteria 42, 43, 44, alongside its further generalisations commonly used to obtain more balanced clusters with minimal cuts 45, 46. Intuitively, the minimum cut partitioning of a network corresponds to the partitioning with the least amount of inter-cluster edges.
Davies-Bouldin index: The DBI quantifies the average similarity between clusters47, calculated based on the distance between clusters relative to the size of the clusters themselves, such that the DBI formula reads:
$\text{DBI}=\frac{1}{M}\sum_{L}^{M}\max_{K\neq L}\left[\frac{\sigma_{K}+\sigma_{L}}{\Delta_{KL}}\right]$ (6)
where $L$ runs over all possible $M$ clusters, $\Delta_{KL}$ denotes the distance between cluster centroids, i.e. $\Delta_{KL}=\left\lVert\mathbf{B}_{K}-\mathbf{B}_{L}\right\rVert$, and $\sigma_{L}$ represents the cluster diameter, calculated as the average distance between the feature vectors and the centroid of the respective cluster:
$\sigma_{L}=\frac{1}{N}\sum_{i}^{N}\left\lVert\mathbf{F}_{i}-\mathbf{B}_{L}\right\rVert$ (7)
where $\mathbf{F}_{i}$ represents the feature vector of node $i$, and the centroid $\mathbf{B}_{L}$ is computed by averaging the features vectors of all $N$ nodes $i$ within cluster $L$. A lower DBI value (bounded from below by 0) indicates better partitioning, as it suggests higher inter-cluster dissimilarity in comparison to intra-cluster similarity. We employed the DBI score as a metric for assessing clustering performance.
@@ 2.2 Markov Chains
The dynamical clustering criterion based on KC we use in this paper is derived from the kinetic properties of an MC. In previous work, we explored the relationship between the MFPTs and KC, providing thorough derivations of all the quantities presented below. Here we briefly summarise the key results required and direct the reader to our previous paper for a detailed reference48.
In a discrete-time, homogeneous MC comprising of $n$ states, the probability vector of a random walk occupying any of the $n$-states at time $t$, $\mathbf{p}(t)$, can be expressed using a stochastic rate matrix, $\mathbf{K}$, and a set of initial occupation probabilities, $\mathbf{p}(0)$, such that:
@@
where $t=l\tau$, since in the discrete case transitions between states occur at integer multiples of a given time interval $\tau$. The quantity $\mathbf{Q}(\tau)=e^{\mathbf{K}\tau}$ is known as the Markov matrix and is commonly used as the fundamental representation of a Markov chain. When using this representation, time propagation corresponds simply to repeatedly multiplying $\mathbf{Q}$ by the current state of the system. To ensure the conservation of probability, we have:
@@
where the index $l$ orders the eigenvalues in descending order. These eigenvalues are upper-bounded by 0 and are guaranteed to be real due to the detailed balanced assumption. The eigenvector corresponding to the largest right eigenvector, $\psi^{(1)}$, is known as the equilibrium probability and is commonly denoted by $\mathbf{p}^{\text{eq}}$, corresponding to the distribution the system converges to as $t\to\infty$. The other eigenvectors contain useful kinetic information corresponding to the relaxation and mixing times of the MC. These eigenvectors can be used to derive an expression for the average time it takes for a random walk starting from state $i$ to first arrive at state $j$, i.e., the MFPTs of an MC, denoted $t_{ji}$:
@@
where the fact that $\sum_{j}\phi^{(l)}_{j}$ $\psi^{(l)}_{j}=1,\forall l$ and $\sum_{j}\psi^{(l)}_{j}=\delta_{l1}$ was used 48. Intuitively, therefore, the Kemeny constant is an important quantity to measure the overall timescales present in a dynamical system, as it is equal to the sum of all relaxation times, $1/|\lambda_{l}|$.
By considering a clustering with the assignment matrix, $\mathbf{S}$, from Eq. (3), we can define the equilibrium probability of cluster $L$, $P^{\text{eq}}_{L}$, as simply the sum of the populations of the individual nodes:
$P^{\text{eq}}_{L}=\sum_{i\in L}p_{i}S_{iL}$ (14)
KC can then be redefined in terms of a partitioned MC’s KC, $\zeta_{\textbf{S}}=\sum_{L}P^{\text{eq}}_{L}\hat{t}_{LK}$, where $\hat{t}_{LK}$ is the coarse-grained mean first passage time from state $K$ to $L$. We can define $\hat{t}_{LK}$ via 48:
$\hat{t}_{LK}=\frac{1}{P_{K}^{\rm eq}P_{L}^{\rm eq}}\sum_{i\in K,j\in L}p^{\rm eq}_{j}p^{\rm eq}_{i}t_{ji}-\frac{1}{(P_{L}^{\rm eq})^{2}}\sum_{i,j\in L}p^{\rm eq}_{j}p^{\rm eq}_{i}t_{ji},$ (15)
which leads to a variational form compared with the unclustered KC29:
$\zeta=\zeta_{\textbf{S}}+\sum_{L}^{M}\frac{1}{P^{\text{eq}}_{L}}\sum_{i,j\in L}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}$ (16)
Rearranging this equation yields our partitioning metric $\Delta K$:
$\zeta-\zeta_{\textbf{S}}=\Delta K=\sum_{L}^{M}\frac{1}{P^{\text{eq}}_{L}}\sum_{i,j\in L}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}=\sum_{L}^{M}\frac{1}{P^{\text{eq}}_{L}}\sum_{i,j}p^{\text{eq}}_{j}p^{\text{eq}}_{i}t_{ji}S_{iL}S_{jL}$ (17)
From this equation, it is clear that for the trivial single-state partitioning $M=1$, $\Delta K=\zeta$ holds, and when $M=N$, the value of $\Delta K$ is 0. As KC represents the sum of relaxation times in a system, and a clustering speeds up relaxation times by removing intra-cluster relaxation times, maximizing the KC of a proposed partitioning aims to preserve the original timescales as closely as possible. From Eq. (17) (with the last equality requiring a crisp clustering), we can see that maximizing the KC is identical to minimizing $\Delta K$, and so this provides a useful metric of how kinetically disconnected the clusters are, in a form suitable for gradient-based optimization methods like neural networks.